# Seeded Topology Optimization for Commercial Foundry Integrated Photonics

### v1

@@
Seeded Topology Optimization for Commercial Foundry Integrated Photonics
@@
journal: opticajournalarticletype: Research Article
{abstract*}
We present a seeded topology optimization methodology for integrated photonic devices fabricated on foundry platforms that yields improved performance compared to traditional topology optimization. We employ blurring filters and a DRC correction algorithm to more readily meet design rule checks yielding devices with fewer artifacts and improved correlation between simulation and measurements. We apply this process to an ultra-compact TE modal multiplexer, a TE mode converter, a polarization rotator, and a high-contrast grating reflector. The measured insertion loss of the TE mode converter was reduced from 1.37 dB to 0.64 dB through this optimization strategy. This approach enables the use of physics-informed device topologies in inverse design and maintains compliance to foundry constraints throughout optimization.
1 Introduction
Inverse design is a rapidly evolving method that is used in the design and optimization of integrated photonic devices to create compact structures with record performance that exploit non-intuitive geometries. Density-based topology optimization (TO) is a flexible form of inverse design that allows each voxel of a design (design parameters) to continuously evolve between two or more materials towards an optimal device topology [1, 2, 3, 4]. In density-based TO, two or more Maxwell simulations are performed each iteration to determine gradients of the design parameters with respect to one or more user-specified figures of merit (FOM) using the adjoint variable method [5, 1]. The gradients are linearly combined to update the design region toward a locally optimum topology.
Devices designed for fabrication at a commercial foundry can be manufactured at high volumes after a successful validation process. This ability to scale is achieved by the requirement that devices conform to stringent design rules checks (DRC) to ensure accurate fabrication with high yield [6]. We can exploit the gradient descent method used in TO by generating gradients that iterate the design to a condition where design rules are met. In our TO pipeline, we have previously implemented algorithms that calculate gradients to meet geometric linewidth constraints (GLC) and area constraints (AC) which are then linearly combined with the FOM gradient each iteration [7].
In traditional TO methodologies, the design parameters are gray-scale i.e., allowed to take on any material value between the two or more materials available on that layer in the material stack. However, to accurately detect DRC violations in a device topology, the device must be sufficiently binary (i.e. every voxel is close to 0 (void) or 1 (solid)) making it challenging to effectively apply DRC constraints. When binarizing the design parameters, a locally optimum topology in the gray-scale phase of the optimization may not translate to a locally optimum binary topology, resulting in a reduction in performance in late stages of TO. This effect is amplified when fabrication-based constraints are included in the optimization, as conflicting gradients (FOM vs. AC/GLC) may cause the optimizer to stall when evolving the geometry to satisfy DRC results in a drop in performance [8, 9]. Hence, there is a need for a modified topology optimization algorithm that maintains device fabricability while exploring the local design space around a functional seed geometry.
In this work we present a seeded TO methodology in which a known functional device geometry, the seed, is iteratively processed and optimized using density-based TO achieving performance beyond the capabilities of traditional TO. A limited blurring filter is applied to enable perturbation of the topology such that the design space around the seed can be explored. We develop a DRC correction algorithm that is catered to the seeded TO process which efficiently resolves foundry constraints while allowing TO to improve device performance. While TO has the potential to discover completely novel geometries, seeded TO focuses efforts to create strictly fabricable devices. We note that the seed can originate from physics-based conventional design strategies. An example is the seeded optimization of high-contrast gratings (HCGs) originally designed using parameter optimization [10, 11]. Shape optimization is a similar inverse design methodology in which a user defines a boundary that is adjusted throughout optimization to maximize performance [12, 13]; however our design methodology allows for changes in the device topology (i.e., elimination and creation of holes/islands) and includes our highly effective DRC correction algorithm. We apply seeded TO to foundry-fabricated inverse designed devices and present an improved design pipeline that utilizes the global optimization features of density-based TO along with the refined, local optimization capability of seeded TO.
2 Topology Optimization Overview
Density-based TO parameterizes the design region such that each voxel can vary between "solid" (high index material) and "void" (low index material) between the two or more materials available on that layer. Typically, every voxel in the design region is initialized to 0.5, meaning the permittivity is halfway between the solid (1) and the void (0) so each voxel can evolve to either. The user specifies the design region lateral dimensions, the location and size of all optical input and output ports (typically waveguides but coupling to a fiber is another example), sources and monitors, and formulates one or more FOMs to minimize. Our unique solver is a hybrid time/frequency-domain adjoint-variable method that readily enables solutions across a wide spectrum via the open-source finite-difference time-domain (FDTD) maxwell solver MEEP [14]. This method allows for the inclusion of multiple FOMs for a single device and enables constraint-based TO [7].
@@ 2 Topology Optimization Overview
In addition to the FOM-based gradients, multiple DRC-based gradients are used to optimize for GLC (minimum linewidth and linespacing) and AC (minimum area and enclosed area) in our TO pipeline. The gradient for GLC is generated using a chosen set of erosions and dilations that identify inflection regions which violate DRC, but can also be generated using morphological transforms [15, 16, 7, 17, 18]. AC includes the minimum area and enclosed area which are implemented using an indicator function that identifies the problematic areas to generate a gradient that encourages the holes/islands to dilate or erode. The minimum curvature radius is another DRC constraint, however the curvature is implicitly enforced by the GLC implementation and the conic filter ($w$) used to map the design parameters [7].
@@
To manage the trade-off between performance and foundry constraints we utilize user-set physical programming bounds that weight the effect of the performance-based FOM and foundry constrains (GLC, and AC) to create the final gradient used in each stage of the optimization [19, 20]. This allows the designer to increase the effect of constraints on the gradient which helps to prevent the optimization from stalling by favoring DRC compliance. These bounds are critical to ensuring the final device is DRC clean but often require hyperparameter tuning through heuristic approaches to mitigate performance loss when AC and GLC physical programming bounds are reduced. Our seeded TO methodology relies less on these bounds, allowing streamlined, efficient optimization that maintains performance with foundry constraints applied.
2.1 Design Parameter Mapping
@@
where $w$ is a conic filter and $\tilde{\rho}$ are the filtered design parameters [21]. The filtered design parameters are then projected using a differentiable, nonlinear function:
@@
where $\beta$ is a threshold parameter gradually increased throughout the optimization process to binarize the device, $\eta$ is a threshold parameter set to 0.5, and $\bar{\rho}$ are the mapped design parameters [22]. The permittivity is interpolated from the mapped design parameters:
@@
where the relative permittivity of each voxel $\epsilon_{r}$ varies between the void ($\epsilon_{min}$) and solid permittivity ($\epsilon_{max}$). This linear interpolation scheme works well for silicon photonic devices; however, nonlinear interpolation schemes may be more suitable for other design problems [23].
3 Seeded Topology Optimization
Figure 1: Overview of the seeded TO process applied to a seed with poor initial performance and significant DRC violations. (a) 16 iterations (M) of seeded TO are applied to optimize transmission to an output waveguide. (b) Overview of one iteration of seeded TO. First, a blurring filter is applied to the binarized design parameters ($\sigma$) to allow the edges of the device to be perturbed during topology optimization. Second, a DRC correction algorithm is applied to the blurred design parameters ($\tilde{\sigma}$), adding material in areas violating minimum linewidth/area and removing material in areas violating minimum line-spacing/enclosed area. Finally, design parameters of TO ($\rho$) are set to the DRC corrected design parameters ($\bar{\sigma}$) and multiple iterations of density-based TO are applied to the device.
The goal of seeded TO is to take a known functional device geometry and improve performance based on a user-specified FOM while maintaining device fabricability. This requires careful processing of the seeded design parameters so the topology can be effectively enhanced using density-based TO. As an extreme example we use seeded TO to optimize a flawed seed with many DRC violations, achieving a fabricable, semi-functional device (Fig. 1a). The DRC violations make this seed ideal for demonstrating how seeded TO can improve device performance while achieving and maintaining DRC clean status. The test cases in this work use functional seeds that comply with DRC and have good initial performance to demonstrate the ability of seeded TO in finding superior local optima.
@@ 3 Seeded Topology Optimization
where $\sigma$ is the binarized design parameters are r is the position vector of a voxel (Fig. 1b). The blurred design parameters ($\tilde{\sigma}$) are then computed:
@@
where $w_{b}$ is a the blurring filter that enables perturbation of the topology during TO. The filter is typically set to a box averaging filter of a user-specified size, however Gaussian filters have also been tested and shown to be effective. The DRC correction algorithm is then applied based on indicator functions used to identify locations in the device topology with DRC violations. The GLC violation indicators (minimum linewidth: $I_{lw}$, minimum linespacing: $I_{ls}$) are found using the open-source software imageruler [24, 25], while the AC indicators (minimum area: $I_{a}$, minimum enclosed area: $I_{ea}$) are determined using a contour detection algorithm [7]. The DRC corrected design parameters ($\bar{\sigma}$) are:
@@
This expands or dilates regions in the mapped topology to algorithmically force DRC compliance beyond inclusion of the constraints in TO. The latent design parameters ($\rho$) of TO are set to the filtered and processed design parameters and 10-20 iterations of TO are performed, allowing the topology to reach a new local minimum. AC and GLC optimization is included in the TO stage, ensuring the final device satisfies DRC. This cycle is repeated until a high-performing, DRC clean device is achieved (Fig. 1a).
3.1 Improved Inverse Design Pipeline
Our new design pipeline begins first with seed generation, this can be accomplished with a variety of tools including traditional TO wherein a device is first optimized with respect to the FOM in traditional TO, followed by optimization with foundry constraints, and completed with seeded TO. This design evolution for a TE modal multiplexer (MMUX) is depicted in Fig. 2. A TE MMUX converts the fundamental mode of 2 single mode waveguides to the TE0 and TE1 modes in a multimode waveguide. In our traditional TO implementation the threshold parameter $\beta$ is typically increased every 20 iterations to slowly binarize the design parameters while the optimizer minimizes the objective function (Fig. 2a) [1]. Before foundry constraints are applied, the FOM often achieves a minimum as the design region is not fully binary (Fig. 2c). AC and GLC optimization is included in the later stages of TO, resulting in reduced performance as the device binarizes and adjusts to satisfy DRC. When constraints are added to TO, the GLC implementation causes the device to binarize but may leave some AC violations (Fig. 2d). To resolve the AC violations, the user set physical programming constraint bounds are further restricted, causing the optimizer to focus on AC over the FOM (Fig. 2e).
Figure 2: (a) TO evolution of a TE modal multiplexer. Initially the design region is set to uniform gray area (b) and optimized solely for the FOM(s) with $\beta$ increasing every 20 iterations to slowly binarize the device. Before the design is fully binarized , constraints are added to the optimization (AC: area constraint, GLC: geometric linewidth constraint) (c). This binarizes the device (d) and the hyperparameters are then tuned to force the device to satisfy DRC (e). Both binary devices ((f) TO binary and (g) TO DRC clean) are used as seeds for seeded TO where the design region is filtered every 12 TO iterations. If no contour detected by the AC algorithm violates DRC, AC is not applied; therefore the AC evaluation does not exist for some iterations.
The seed topology is not required to be DRC clean, but it must be binary. Seeded TO was performed on the MMUX 30 iterations after constraints are applied (Fig. 2f) and after TO achieves a DRC clean device (Fig. 2g) [9]. The FOM for both cases reduces throughout the optimization, however the AC and GLC evaluation was much lower for the MMUX seed that was DRC clean. The resulting MMUXs are almost identical, both satisfying DRC with similar performance. This allows the user to end TO early and achieve an optimal device topology, reducing the computational cost of an optimization.
3.2 Blurring Filter Study

### v5

@@
Abstract
We present a seeded topology optimization methodology for integrated photonic devices fabricated on foundry platforms that yields improved performance compared to traditional topology optimization. We employ blurring filters and a design rule check correction algorithm to more readily meet fabrication constraints, resulting in devices with fewer artifacts and improved correlation between simulation and measurements. A statistical study is performed on a 2D modal multiplexer, revealing that 87% of devices optimized using this strategy conform to foundry constraints, compared to 13% of devices optimized using traditional TO. We apply seeded topology optimization to an ultra-compact TE modal multiplexer, a TE mode converter, a polarization rotator, and a high-contrast grating reflector. Using this optimization strategy, the measured insertion loss of the TE mode converter was reduced from < 1.50 dB to < 0.64 dB, and the measured TE1 insertion loss of the TE modal multiplexer was reduced from < 3.95 dB to < 1.38 dB over C-band. This approach enables a two-step inverse design process, merging of physics-informed design strategies with inverse design, and ensures strict compliance with foundry constraints throughout optimization.
preprint: AIP/123-QED
I Introduction
Inverse design is a rapidly evolving method that is used in the design and optimization of integrated photonic devices to create compact structures with record performance that exploit non-intuitive geometries. Density-based topology optimization (TO) is a flexible form of inverse design that allows each voxel of a design (the design parameters) to continuously evolve between two or more materials towards an optimal device topology based on user defined constraintsHammond et al. (2022a); Molesky et al. (2018); Su et al. (2018); Christiansen and Sigmund (2021); Sell et al. (2017). In density-based TO, two or more Maxwell simulations are performed each iteration to determine the gradient(s) of one or more user-specified figures of merit (FOM) with respect to all design parameters using the adjoint variable method Strang (2007); Hammond et al. (2022a). The gradients for each FOM are linearly combined to update the design region toward a locally optimum topology.
Devices designed for fabrication at a commercial foundry are readily manufactured at high volume after a successful validation process. This ability to scale is achieved by the requirement that devices conform to stringent design rules checks (DRC) to ensure accurate fabrication with high yield Liehr et al. (2020). We can exploit the gradient descent method used in TO by generating gradients that iterate the design to a condition where design rules are met. In our traditional TO pipeline, we have previously implemented algorithms that calculate gradients to meet geometric linewidth constraints (GLC) and area constraints (AC), which are then linearly combined with the FOM gradient each iteration Hammond et al. (2021).
In traditional TO methodologies, the design parameters are initially gray-scale, i.e., allowed to take on any material value between the two or more materials available on that layer in the material stack. However, to accurately detect DRC violations, the device must be sufficiently binary (i.e., every voxel is close to 0 (void) or 1 (solid)), making it challenging to effectively apply DRC constraints. When binarizing the design parameters, a high-performing local optimum in the gray-scale phase of the optimization may not translate to a high-performing local optimum binary topology. Gradient-based optimizers are known to get trapped in local minima valleys or saddle points, which limits the performance achievable through inverse design Yeung et al. (2022); Zhou et al. (2015); Ballew et al. (2023). This effect is amplified when fabrication-based constraints are included in the optimization, as conflicting objectives and constraints (FOM vs. AC/GLC) may cause the optimizer to stall when evolving the geometry to satisfy DRC results in a drop in performance Pesch et al. ; Hiesener et al. . Hence, there is a need for a modified topology optimization algorithm that maintains device fabricability while exploring the local design space around a functional seed geometry.
In this work, we present a seeded TO methodology in which a known functional device geometry, the seed, is iteratively processed and optimized using density-based TO. We demonstrate improved performance compared to traditional TO or conventionally designed seed devices, all designed for a commercial foundry. A limited blurring filter is applied to enable perturbation of the topology such that the design space around the seed can be explored. We develop a DRC correction algorithm that is catered to the seeded TO process, which efficiently resolves foundry constraints while allowing a nested TO stage to improve device performance.
While traditional TO has the potential to discover completely novel geometries, seeded TO focuses efforts to create strictly fabricable devices. We note that the seed can originate from physics-based conventional design strategies. One example of this is the seeded optimization of high-contrast gratings (HCGs) originally designed using parameter optimization Hiesener, Pesch, and Ralph (2024); Wong et al. . Shape optimization is a similar inverse design methodology in which a user defines a boundary that is adjusted throughout optimization to maximize performance Lalau-Keraly et al. (2013). Previous work has explored seeding shape optimization with a physics-informed initial structure for multimode interferometer or metasurface design Michaels, Wu, and Yablonovitch (2019); Dainese et al. (2024). While shape optimization restricts the device topology, allowing for strict enforcement of DRC, our design methodology enables changes in the device topology (i.e., elimination and creation of holes/islands) and includes a DRC correction algorithm that provides more control over minimum linewidths and areas throughout the topology. We apply seeded TO to foundry-fabricated inverse-designed devices and present an improved design pipeline that utilizes the coarse global optimization features of density-based TO followed by the fine, local optimization capability of seeded TO.
II Traditional Topology Optimization Overview
Density-based TO parameterizes the design region such that each voxel can vary between "solid" (high index material) and "void" (low index material) between the two or more materials available on that layer. Typically, every voxel in the design region is initialized to 0.5 or random noise between 0 and 1, meaning the permittivity is between the solid (1) and the void (0), so each voxel can evolve to either. The user specifies the design region lateral dimensions, the location and size of all optical input and output ports (typically waveguides butt-coupling to a fiber is another example), sources and monitors, and formulates one or more FOMs to minimize. Our unique solver is a hybrid time/frequency-domain adjoint-variable method that readily enables solutions across a wide spectrum via the open-source finite-difference time-domain (FDTD) Maxwell solver MEEP Oskooi et al. (2010). This method allows for the inclusion of multiple FOMs for a single device and enables constraint-based TO Hammond et al. (2021). We formulate TO as a multi-objective minimization problem with $N$ objective functions $(f_{1},...,f_{N})$ subject to Maxwell’s equations at $M$ frequency points, bounds for the design parameters $(\rho)$, and $K$ constraint functions $(g_{1},...,g_{K})$:
@@ II Traditional Topology Optimization Overview
where $\bar{g}_{i},\ i\in\{1,...,N+K\}$ are differentiable spline-based scaling functions applied to both the objectives and the constraints that are generated using a user defined list of bounds referred to as physical programming bounds ($\mathbf{q}$) Messac (1996); Probst et al. (2024). The returns of each objective/constraint are mapped such that they are optimized on a unified scale, giving the designer more control over the effect of the objective functions and the constraints throughout optimization.
The adjoint variable method is implemented using the in-built solver in MEEP to compute the gradient of the FOM with respect to the design parameters Hammond et al. (2022a); Oskooi et al. (2010). The gradient is then backpropagated through the parameterization using a vector-Jacobian product implemented via the open-source software package Autograd Maclaurin, Duvenaud, and Adams (2015). The latent design variables are optimized using the globally convergent method of moving asymptotes (GCMMA) provided by the open-source nonlinear optimization package NLopt Johnson . The GCMMA optimizer produces a sequence of iteration points that are guaranteed to converge to a set of Karush-Kuhn-Tucker points; however, the optimizer may converge to a local minimum with poor performance due to conflicting objectives or constraints Svanberg (2002). A different optimizer may reduce the chance that the optimization gets trapped in a local minimum; however, the GCMMA optimizer allows for the large number of design parameters and constraints required for commercial foundry applications. A comparison of common optimizers used in TO is given in Supp. A.
@@
Though the lithography resolution may limit maximum device performance, traditional TO has been successfully demonstrated to produce ultra-compact, high-performing devices in the constrained design space required for commercial foundries Zhou et al. (2015); Hammond et al. (2021); Piggott et al. (2020); Khurana et al. (2023); Slaby, Hammond, and Ralph ; Hammond et al. (2022b). The constraints applied to optimizations in this work are GLC (minimum linewidth and linespacing) and AC (minimum area and enclosed area), constraints imposed by the commercial foundry. The gradient for GLC is generated using a chosen set of erosions and dilations that identify inflection regions which violate the minimum linewidth or linespacing, but can also be generated using morphological transforms Vercruysse et al. (2019); Zhou et al. (2015); Hammond et al. (2021); Piggott et al. (2020, 2017). AC includes the minimum area and enclosed area, which are implemented using an indicator function that identifies the violating areas to generate a gradient that encourages the holes/islands to dilate or erode. The minimum radius of curvature is another DRC constraint; however, the curvature is implicitly enforced by the GLC implementation and the conic filter ($w$) used to map the design parameters Hammond et al. (2021). See Supp. B for a mathematical description of the GLC and AC implementations.
@@
The user-set physical programming bounds allow the designer to strategically increase the effect of constraints on the gradient, ensuring that the final device satisfies DRC. These bounds are critical to ensuring the final device is DRC clean but often require hyperparameter tuning through heuristic approaches to mitigate performance loss when AC and GLC physical programming bounds are reduced. Our seeded TO methodology relies less on these bounds, allowing streamlined, efficient optimization that maintains performance with foundry constraints applied.
II.1 Design Parameter Mapping
@@
where $w$ is a conic filter and $\tilde{\rho}$ are the filtered design parameters Sigmund and Maute (2013). The filtered design parameters are then projected using a differentiable, nonlinear function:
@@
where $\beta$ is a threshold parameter gradually increased throughout the optimization process to binarize the device, $\eta$ is a threshold parameter set to 0.5, and $\bar{\rho}$ are the mapped design parameters Wang, Lazarov, and Sigmund (2011). The permittivity is interpolated from the mapped design parameters:
@@
where the relative permittivity of each voxel $\epsilon_{r}$ varies between the void ($\epsilon_{min}$) and solid permittivity ($\epsilon_{max}$). This linear interpolation scheme works well for silicon photonic devices; however, nonlinear interpolation schemes may be more suitable for other design problems Christiansen et al. (2019).
III Seeded Topology Optimization
Figure 1: Overview of the seeded TO process applied to a seed with poor initial performance and significant DRC violations. (a) 16 iterations (M) of seeded TO are applied to optimize transmission to an output waveguide. (b) Overview of one iteration of seeded TO. First, a blurring filter is applied to the binarized design parameters ($\sigma$) to allow the edges of the device to be perturbed during topology optimization. Second, a DRC correction algorithm is applied to the blurred design parameters ($\tilde{\sigma}$), adding material in areas violating minimum linewidth/area and removing material in areas violating minimum line-spacing/enclosed area. Finally, design parameters of TO ($\rho$) are set to the DRC corrected design parameters ($\bar{\sigma}$) and multiple iterations of nested TO are applied to the device.
The goal of seeded TO is to take a known functional device geometry and improve performance based on a user-specified FOM while maintaining device fabricability. This requires careful processing of the seeded design parameters so the topology can be effectively enhanced using density-based TO. The "seeded" aspect of this methodology refers to initializing the design parameters of seeded TO to a functional device topology. As an extreme example, we use seeded TO to optimize a flawed seed with many DRC violations, achieving a fabricable, semi-functional device (Fig. 1a). The DRC violations make this seed ideal for demonstrating how seeded TO can improve device performance while achieving and maintaining DRC clean status. The test cases in this work use functional seeds that comply with DRC and have good initial performance to demonstrate the ability of seeded TO in finding superior local optima.
@@ III Seeded Topology Optimization
where $\sigma$ is the binarized design parameters and r is the position vector of a voxel (Fig. 1b). The variable $\sigma$ is used to demarcate the seeded TO design parameters as they are processed differently than the design parameters ($\rho$) used in traditional TO. This binarization stage gives the designer control over the level of blur applied to the design parameters, as any blur introduced via TO or previous seeded TO iterations is removed. The blurred design parameters ($\tilde{\sigma}$) are then computed:
@@
where $w_{b}$ is the blurring filter. This filter is distinct from the conic filter ($w$) used described in sec. 2.1, which enables a smoothed projection of the design parameters to the simulated permittivities, whereas the blurring filter enables perturbation of the topology by the nested TO stage of seeded TO. This blurring stage promotes improvement via TO, since TO can more effectively perturb gray-scale design parameters compared to binary design parameters. The blurring filter is typically set to a box averaging filter of a user-specified size; however, Gaussian filters have also been tested and shown to be effective. The DRC correction algorithm is then applied based on indicator functions used to identify locations in the device topology with DRC violations. The GLC violation indicators (minimum linewidth: $I_{lw}$, minimum linespacing: $I_{ls}$) are found using the open-source software imageruler Chen et al. (2024); Oskooi , while the AC indicators (minimum area: $I_{a}$, minimum enclosed area: $I_{ea}$) are determined using a contour detection algorithm Hammond et al. (2021). All the indicators are calculated using the binarized design parameters $\sigma$. The DRC corrected design parameters ($\bar{\sigma}$) are:
@@
This expands or dilates regions in the mapped topology to algorithmically force DRC compliance beyond inclusion of DRC constraints in TO. Fewer violating inflection regions and contours exist after the DRC correction algorithm is applied, reducing the evaluation of the constraints, which in turn reduces the magnitude of the constraint gradients generated in the nested TO stage. The latent design parameters ($\rho$) of TO are set to the filtered and processed design parameters ($\bar{\sigma}$):
$\rho=\bar{\sigma}$ (8)
and 10-20 iterations of TO are performed (referred to as nested TO). Sufficient nested TO iterations must be used to ensure the device reaches a new local minimum, but not so many that the performance plateaus. Nested TO is nearly identical to traditional TO, using the same design parameter mapping scheme as described in Section 2.1 with a high threshold parameter ($\beta$) value (at or above the $\beta$ value used in the final iterations of traditional TO when using TO to generate the seed). AC and GLC must be included in nested TO to mitigate any opposition between the DRC correction and nested TO stages of seeded TO. If DRC constraints are not included in nested TO, nested TO may resist the DRC correction step, causing oscillatory behavior that prevents convergence. This seeded TO cycle is repeated until a high-performing, DRC clean device is achieved.
In traditional TO, careful hyperparameter tuning is required to gradually optimize the device towards a binary topology that satisfies constraints. The traditional TO design pipeline can be transformed into a two-step design process, first with an initial coarse optimization using traditional TO, followed by a fine optimization using seeded TO. This two-step design process uses the broad search space of traditional TO and the DRC correction capability of seeded TO for efficient device optimization.
III.1 2D Example: Modal Multiplexer
Figure 2: (a) Optimization process flow for a modal multiplexer using either traditional TO (red) or seeded TO (gold). The design parameters for each run are initialized to randomly generated noise. The threshold parameter ($\beta$) is gradually increased every 24 iterations, and constraints are optimized for starting on iteration 48. 100 independent runs were performed for both traditional TO and seeded TO. (b) Standard swarm plot of the average transmission of both modes over the optimization band for all 100 devices optimized using each method. Pie charts indicate the number of devices that pass DRC (blue), fail geometric linewidth constraints (GLC, orange), fail area constraints (AC, purple), and fail both constraints (red).
To demonstrate the utility of seeded TO, we perform a statistical study on a modal multiplexer (MMUX), where we design using both a traditional TO and a seeded TO pipeline. We begin by initializing each voxel of the design parameters to a random number from a uniform distribution over the interval $(0,1)$ and optimizing solely for the user-defined FOM while gradually increasing the threshold parameter ($\beta$) to binarize the device. Once the device is sufficiently binary, foundry constraints are added to the optimization. In the traditional TO pipeline, the area and linewidth constraints are increasingly enforced by reducing the physical programming bounds while increasing the threshold parameter ($\beta$) to binarize the device. Reducing the physical programming bounds scales the magnitude of the constraint functions compared to the objective functions, encouraging the optimizer to prioritize minimizing the constraint evaluations. In the seeded TO pipeline, seeded TO is applied to the device shortly after constraints are added to the optimization.
100 MMUX optimizations were performed using 2D FDTD simulations for each design pipeline (traditional TO and seeded TO) (Fig. 2a). The design parameters of each optimization were initialized with distinct randomly generated noise; however, TO and seeded TO were performed using the same set of hyperparameters for all runs. This MMUX converts the lowest two TE modes in a multimode waveguide to the fundamental mode of two single-mode waveguides (Sec. 4.1). This optimization problem tends to generate a hole near the multimode waveguide interface which is limited by the minimum area and linespacing constraints, therefore it is an excellent test of the capability of seeded TO in strictly enforcing DRC. Typically, when performing traditional TO, the designer updates the hyperparameters dynamically throughout the optimization to efficiently resolve foundry constraint violations; however, this is challenging when optimizing for multiple FOMs (both modes) due to counteracting gradients. Seeded TO uses both the DRC correction algorithm and nested TO constraint optimization to resolve DRC violations; therefore, it requires minimal hyperparameter tuning to achieve a working device.
The resulting devices were binarized, checked for compliance with GLC and AC, and simulated (Fig. 2b). 68 out of 100 seeded TO optimizations yielded devices with superior performance than their traditional TO counterparts. Additionally, 87 of the seeded TO devices conformed to all DRC constraints, while only 13 of the traditional TO devices passed DRC.
Many of the traditional TO devices failed only AC or GLC; these topologies could be adjusted manually after optimization to conform to DRC with minimal performance degradation Hiesener et al. . The highest-performing DRC clean devices optimized using traditional TO achieve the same performance as the highest-performing seeded TO devices, indicating that the selected hyperparameters are effective for these runs, but not for most other traditional TO runs. A majority of the seeded TO devices have above 80% transmission while maintaining DRC compliance, indicating that the seeded TO process is less sensitive to hyperparameters compared to traditional TO. Though both may satisfy foundry constraints, the seeded TO devices are more suitable for fabrication compared to the traditional TO devices due to the smooth topologies that are generated in the optimization process (Sec. 4).
III.2 Blurring Filter Study
Figure 3: (a) Traditional TO was used to design a $\upmum ×\times 2.42.4 μ\upmum$ modal multiplexer in a 2D simulation environment which contains DRC violations. In seeded TO the design parameters are blurred with a DRC correction algorithm applied. Identical to traditional TO, the parameters are passed through a mapping function to return the permittivity values to be used in the Maxwell simulation. (b) Comparison of varying box filter dimensions used for the seeded TO blurring filter, with DRC violations highlighted.